# Metrics on completely positive maps via noncommutative geometry

### v1

@@ 4.1 The map $\omega_{\tau}$ and trace channels
${A\odot B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F\odot\operatorname{id}_{B^{\rm op}}}$${B\odot B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{\tau}}$${\mathbb{C}.}$
We will often make use of this fact.
We are furthermore interested in how $\omega_{\tau}$ interacts with different $C^{*}$-norms on $A\odot B^{\rm op}$. In particular, if $\delta$ is a $C^{*}$-norm on $A\odot B^{\rm op}$, we want to now for which linear maps $F\colon A\to B$ do we have that $\omega_{\tau}(F)$ is continuous with respect to $\delta$. To that end, we give the following definition.
@@
Let $A$ and $B$ be $C^{*}$-algebras, and suppose that $\tau$ is a trace on $B$. For a $C^{*}$-norm $\delta$ on $A\odot B^{\rm op}$ define the set $\mathcal{L}_{\tau}^{\delta}(A,B)$ to be the set of linear maps $F\colon A\to B$ such that $\omega_{\tau}(F)$ is continuous with respect to $\delta$. In this case we use the notation $\omega_{\tau}^{\delta}(F)$ to denote the (unique) extension of $\omega_{\tau}(F)$ to $A\otimes_{\delta}B^{\rm op}$. If either $A$ or $B$ is nuclear, no extra decoration will be made to denote the extension of $\omega_{\tau}(F)$ to $A\otimes B^{\rm op}$.
In the special cases where $\delta=\|\cdot\|_{\rm min}$ and $\delta=\|\cdot\|_{\rm max}$ we will simply use the notation $\mathcal{L}_{\tau}^{\rm min}(A,B)$ and $\mathcal{L}_{\tau}^{\rm max}(A,B)$ respectively.
@@
In this definition we are considering the possibility that $\omega_{\tau}(F)$ may be continuous in some $C^{*}$-norm on $A\odot B$ without assuming anything about the continuity of $\mu_{\tau}$. If $\mu_{\tau}$ is continuous (as it always is with respect to the maximal norm), then continuity of $\omega_{\tau}(F)$ follows by continuity of $F\odot\operatorname{id}_{B^{\rm op}}$.
@@
We will show that the completely positive maps from $A$ to $B$ are precisely the maps $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$ for which $\omega_{\tau}^{\rm max}(F)$ is positive. First we show that if $F\in\operatorname{CP}(A,B)$, then $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$.
@@
Whenever $F\in{\rm Dec(A,B)}:={\rm Span}(\operatorname{CP}(A,B))$ we have that $\omega_{\tau}(F)$ is continuous in the maximal norm on $A\odot B^{\rm op}$. This follows in the same way as in the proof of Proposition 4.6, now using [Pis20a, Proposition 6.11], which implies that
@@
for every $F\in{\rm Dec(A,B)}$. On the other hand, if we assume that $\tau$ is amenable, then we may use [Pis20a, Proposition 1.11] to deduce that $\omega_{\tau}(F)$ is continuous in the spatial norm on $A\odot B^{\rm op}$ for every $F\in\operatorname{CB}(A,B)$, where $\operatorname{CB}(A,B)$ denotes the set of all completely bounded maps from $A$ to $B$.
@@
Let $A$ and $C$ be $C^{*}$-algebras, and $B$ and $D$ be $C^{*}$-algebras with traces $\tau_{B}$ and $\tau_{D}$ respectively. Then for any $F\in\operatorname{CP}(A,B)$ and $G\in\operatorname{CP}(C,D)$ we have that
$(\Sigma_{[23]})^{*}\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}(F\otimes G)=\omega_{\tau_{B}}^{\rm max}(F)\otimes\omega_{\tau_{D}}^{\rm max}(G).$ (4.2)
@@
$(\Sigma_{[23]})^{*}\omega_{\tau_{B}\otimes\tau_{D}}^{\rm min}(F\otimes G)=\omega_{\tau_{B}}^{\rm min}(F)\otimes\omega_{\tau_{D}}^{\rm min}(G).$ (4.3)
@@
By Lemma 2.7 and Proposition 4.6 we have that the all the maps in the following diagram are well-defined.
${\operatorname{CP}(A,B)\times\operatorname{CP}(C,D)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\otimes}$$\scriptstyle{\omega_{\tau_{B}}^{\rm max}\times\omega_{\tau_{D}}^{\rm max}}$${(A\otimes_{\rm max}B^{\rm op})^{*}\times(C\otimes_{\rm max}D^{\rm op})^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\otimes}$${\operatorname{CP}_{\tau_{B}\otimes\tau_{D}}(A\otimes_{\rm max}C,B\otimes_{\rm max}D)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}}$${(A\otimes_{\rm max}C\otimes_{\rm max}B^{\rm op}\otimes_{\rm max}D^{\rm op})^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\Sigma_{[23]})^{*}}$${(A\otimes_{\rm max}B^{\rm op}\otimes_{\rm max}C\otimes_{\rm max}D^{\rm op})^{*}}$ (4.4)
We have that this diagram commutes if and only if Equation 4.2 hold for all $F\in\operatorname{CP}(A,B)$ and $G\in\operatorname{CP}(C,D)$. Hence, we simply want to show that this diagram commutes. For $a\in A$, $b\in B$, $c\in C$, and $d\in D$ we have that
@@
Now suppose that $B$ and $D$ are unital, and that $\tau_{B}$ and $\tau_{D}$ are amenable. By [BO08, Proposition 6.3.5] we get that $\tau_{B}\otimes\tau_{D}$ is an amenable trace on $B\otimes_{\rm min}D$. Hence, everything appearing in the diagram presented in Equation 4.4 is still well-defined with min in place of max by Lemma 2.7 and Proposition 4.6. That Equation 4.3 holds then follows by the same computation we did earlier in this proof. ∎
@@
Since the *-isomorphism $\Sigma_{[23]}$ is idempotent, we have that Equation 4.2 and Equation 4.3 give rise to analogous equations
$\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}(F\otimes G)=(\Sigma_{[23]})^{*}(\omega_{\tau_{B}}^{\rm max}(F)\otimes\omega_{\tau_{D}}^{\rm max}(G)),$ (4.5)
@@
$\omega_{\tau_{B}\otimes\tau_{D}}^{\rm min}(F\otimes G)=(\Sigma_{[23]})^{*}(\omega_{\tau_{B}}^{\rm min}(F)\otimes\omega_{\tau_{D}}^{\rm min}(G)).$ (4.6)
We will also have use for these equations.
@@
Let $A$ and $B$ be a $C^{*}$-algebra, and let $\tau$ be a faithful trace on $B$. For any $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$ we have that $F$ is completely positive if and only if $\omega_{\tau}^{\rm max}(F)$ is positive.
@@
Suppose first that $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$ is completely positive. Since $\omega_{\tau}^{\rm max}(F)$ and $\mu_{\tau}^{\rm max}\circ(F\otimes\operatorname{id}_{B^{\rm op}})$ agree on the dense subspace $A\odot B^{\rm op}$ we have that $\omega_{\tau}^{\rm max}(F)$ is given by the following composition:
@@
$(\operatorname{Tr}_{n}\otimes\tau)((\operatorname{id}_{n}\otimes F)(X)Y)$ $=(\omega_{\operatorname{Tr}_{n}\otimes\tau}(\operatorname{id}_{n}\otimes F))(X\otimes Y^{\rm op})$ by Equation 4.1
@@
$=((\Sigma_{[23]})^{*}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(F)))(X\otimes Y^{\rm op})$ by Equation 4.5
@@
We note that if $B$ is unital and $\tau_{B}$ is a tracial state, then $\operatorname{UCP}(A,B)$ is contained in $\operatorname{TC}_{\tau}(A,B)$.
@@
Let $A$ and $B$ be $C^{*}$-algebras with faithful traces $\tau_{A}$ and $\tau_{B}$ respectively. We say that a linear map $F\colon A\to B$ is $(\tau_{A},\tau_{B})$-adjointable if there exists a linear map $F^{\sharp}\colon B\to A$ such that
$\tau_{B}(F(a)b)=\tau_{A}(aF^{\sharp}(b))$ (4.7)
@@
Let $A$ and $B$ be unital $C^{*}$-algebras with faithful traces $\tau_{A}$ and $\tau_{B}$ respectively. If $F\in\operatorname{TC}_{\tau_{B}}(A,B)$ is $(\tau_{A},\tau_{B})$-adjointable, then $F^{\sharp}\in\operatorname{TC}_{\tau_{A}}(B,A)$, and we furthermore have that
$\omega_{\tau_{A}}(F^{\sharp})=\omega_{\tau_{B}}(F)\circ\Sigma^{\rm op}.$ (4.8)
@@
4.2 Relation to the Choi-Jamiolkowski isomorphisms
@@
Let $n,m\in\mathbb{N}$, $A=M_{n}(\mathbb{C})$, $B=M_{m}(\mathbb{C})$, and $\tau$ be the canonical trace on $B$. Let $F\colon A\to B$ be a linear map and denote the standard matrix units in $M_{n}(\mathbb{C})$ by $e_{i,j}$ for $i,j=1,\ldots,n$. We recall, see for instance [Stø13, Chapter 4], that the Choi matrix $C_{F}$ of $F$ is given by
$C_{F}:=\sum_{i,j=1}^{n}e_{i,j}\otimes F(e_{i,j})\,\in\,M_{n}(\mathbb{C})\otimes M_{m}(\mathbb{C})\simeq M_{nm}(\mathbb{C}).$
The assignment $F\mapsto C_{F}$ establishes an isomorphism between $\operatorname{Hom}(A,B)$ and $A\otimes B$. This isomorphism is called the Choi-Jamiolkowski isomorphism.
@@
for all $a\in M_{n}(\mathbb{C})$, $b\in M_{m}(\mathbb{C})$, and that $\widehat{F}$ is positive if (and only if) $F$ is completely positive, in which case the positive density matrix corresponding to $\widehat{F}$ is the transpose of $C_{F}$, cf. [Stø13, Lemma 4.2.3 and Theorem 4.2.7]. Now, let $t$ denote the isomorphism from $B^{\rm op}=M_{m}(\mathbb{C})^{\rm op}$ onto $B=M_{m}(\mathbb{C})$ given by $t(b^{\rm op})=b^{t}$ for all $b\in B$. Then one readily verifies that
@@
Thus, if $F$ is completely positive and we identify $A\otimes B^{\rm op}$ with $A\otimes B$ via the isomorphism $\operatorname{id}_{A}\otimes t$, we get that the density matrix corresponding to the positive functional $\omega_{\tau}(F)$ is the positive matrix $(C_{F})^{t}$.
@@
$\widehat{F}(a\otimes e_{i,j})=F(a)_{i,j}\quad\text{ for every $a\in A$ and $i,j\in\{1,\ldots,m\}$},$
where the $e_{i,j}$’s denote the standard matrix units in $M_{m}(\mathbb{C})$ and $F(a)_{i,j}$ denotes the $(i,j)$-coefficient of the matrix $F(a)$. Then, as shown in [BO08, Proposition 1.5.14], the map $F$ is completely positive if and only if $\widehat{F}$ is positive, and the map $F\mapsto\widehat{F}$ gives a bijective correspondence between $\operatorname{CP}(A,M_{m}(\mathbb{C}))$ and the cone of all positive linear functionals on $A\otimes M_{m}(\mathbb{C})$. Now, it is easy to verify that $\widehat{F}(a\otimes b)={\rm Tr}_{m}(F(a)b^{t})$ for all $a\in A$ and $b\in M_{m}(\mathbb{C})$. Thus, in analogy with Example 4.18, $\widehat{F}$ might be called the dual functional of $F$ on $A\otimes M_{m}(\mathbb{C})$. Moreover, setting $t(b^{\rm op}):=b^{t}$ for all $b\in M_{m}(\mathbb{C})$, it is clear that we have $\omega_{\tau}(F)=\widehat{F}\circ({\rm id}_{A}\otimes t)$.
@@
In a von Neumann algebraic setting, there exist some infinite-dimensional versions of the Choi-Jamiolkowski isomorphism in the literature. See for example [Haa21] and [HKS24] and references therein.
5 Induced metrics on trace channels
As discussed in Section 3, we will in this section determine some mild criteria, under which we can guarantee that stability and chaining hold. All results in this section are stated and proved using seminorms (rather than Lipschitz seminorms). However, the examples we have in mind arise from spectral triples through constructions explained in Section 2.2.
5.1 Stability

### v2

@@ 5.1. The map $\omega_{\tau}$ and trace channels
${A\odot B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F\odot\operatorname{id}_{B^{\rm op}}}$${B\odot B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{\tau}}$${\mathbb{C},}$
(where $\mu_{\tau}$ is defined in section 2.1). We will often make use of this fact.
We are furthermore interested in how $\omega_{\tau}$ interacts with different $C^{*}$-norms on $A\odot B^{\rm op}$. In particular, if $\delta$ is a $C^{*}$-norm on $A\odot B^{\rm op}$, we want to know for which linear maps $F\colon A\to B$ do we have that $\omega_{\tau}(F)$ is continuous with respect to $\delta$. To that end, we make the following definition.
@@
Let $A$ and $B$ be $C^{*}$-algebras, and suppose that $\tau$ is a trace on $B$. For a $C^{*}$-norm $\delta$ on $A\odot B^{\rm op}$ define $\mathcal{L}_{\tau}^{\delta}(A,B)$ to be the set of linear maps $F\colon A\to B$ such that $\omega_{\tau}(F)$ is continuous with respect to $\delta$. In this case we use the notation $\omega_{\tau}^{\delta}(F)$ to denote the (unique) extension of $\omega_{\tau}(F)$ to $A\otimes_{\delta}B^{\rm op}$. If either $A$ or $B$ is nuclear, no extra decoration will be made to denote the extension of $\omega_{\tau}(F)$ to $A\otimes B^{\rm op}$.
In the special cases where $\delta=\|\cdot\|_{\rm min}$ and $\delta=\|\cdot\|_{\rm max}$, we will simply use the notation $\mathcal{L}_{\tau}^{\rm min}(A,B)$ and $\mathcal{L}_{\tau}^{\rm max}(A,B)$, respectively.
@@
In this definition we are considering the possibility that $\omega_{\tau}(F)$ may be continuous in some $C^{*}$-norm on $A\odot B^{\rm op}$ without assuming anything about the continuity of $\mu_{\tau}$. If $\mu_{\tau}$ is continuous (as it always is with respect to the maximal norm), then continuity of $\omega_{\tau}(F)$ follows by continuity of $F\odot\operatorname{id}_{B^{\rm op}}$.
@@
We will show that the completely positive maps $\operatorname{CP}(A,B)$ from $A$ to $B$ are precisely the maps $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$ for which $\omega_{\tau}^{\rm max}(F)$ is positive. First we show that if $F\in\operatorname{CP}(A,B)$, then $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$.
@@
Whenever $F\in{\rm Span}(\operatorname{CP}(A,B))$, i.e., $F$ is decomposable, we have that $\omega_{\tau}(F)$ is continuous in the maximal norm on $A\odot B^{\rm op}$. This follows in the same way as in the proof of proposition 5.6, now using [Pis20a, Proposition 6.11], which implies that
@@
for every $F\in{\rm Span}(\operatorname{CP}(A,B))$. On the other hand, if we assume that $\tau$ is amenable, then we may use [Pis20a, Proposition 1.11] to deduce that $\omega_{\tau}(F)$ is continuous in the spatial norm on $A\odot B^{\rm op}$ for every $F\in\operatorname{CB}(A,B)$, where $\operatorname{CB}(A,B)$ denotes the set of all completely bounded maps from $A$ to $B$.
@@
Let $A$ and $C$ be $C^{*}$-algebras, and $B$ and $D$ be $C^{*}$-algebras with traces $\tau_{B}$ and $\tau_{D}$, respectively. Then for any $F\in\operatorname{CP}(A,B)$ and $G\in\operatorname{CP}(C,D)$ we have that
(5.2) $(\Sigma_{[23]})^{*}\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}(F\otimes G)=\omega_{\tau_{B}}^{\rm max}(F)\otimes\omega_{\tau_{D}}^{\rm max}(G).$
@@
(5.3) $(\Sigma_{[23]})^{*}\omega_{\tau_{B}\otimes\tau_{D}}^{\rm min}(F\otimes G)=\omega_{\tau_{B}}^{\rm min}(F)\otimes\omega_{\tau_{D}}^{\rm min}(G).$
@@
By lemma 2.7 and proposition 5.6 we have that all the maps in the following diagram are well-defined.
(5.4) ${\operatorname{CP}(A,B)\times\operatorname{CP}(C,D)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\otimes}$$\scriptstyle{\omega_{\tau_{B}}^{\rm max}\times\omega_{\tau_{D}}^{\rm max}}$${(A\otimes_{\rm max}B^{\rm op})^{*}\times(C\otimes_{\rm max}D^{\rm op})^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\otimes}$${\operatorname{CP}_{\tau_{B}\otimes\tau_{D}}(A\otimes_{\rm max}C,B\otimes_{\rm max}D)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}}$${(A\otimes_{\rm max}C\otimes_{\rm max}B^{\rm op}\otimes_{\rm max}D^{\rm op})^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\Sigma_{[23]})^{*}}$${(A\otimes_{\rm max}B^{\rm op}\otimes_{\rm max}C\otimes_{\rm max}D^{\rm op})^{*}}$
We have that this diagram commutes if and only if (5.2) hold for all $F\in\operatorname{CP}(A,B)$ and $G\in\operatorname{CP}(C,D)$. Hence, we simply want to show that this diagram commutes. For $a\in A$, $b\in B$, $c\in C$, and $d\in D$ we have that
@@
Now suppose that $B$ and $D$ are unital, and that $\tau_{B}$ and $\tau_{D}$ are amenable. By [BO08, Proposition 6.3.5] we get that $\tau_{B}\otimes\tau_{D}$ is an amenable trace on $B\otimes_{\rm min}D$. Hence, everything appearing in the diagram presented in (5.4) is still well-defined with min in place of max by lemma 2.7 and proposition 5.6. That (5.3) holds then follows by the same computation we did earlier in this proof. ∎
@@
Since the *-isomorphism $\Sigma_{[23]}$ is idempotent, we have that (5.2) and (5.3) give rise to analogous equations
(5.5) $\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}(F\otimes G)=(\Sigma_{[23]})^{*}(\omega_{\tau_{B}}^{\rm max}(F)\otimes\omega_{\tau_{D}}^{\rm max}(G)),$
@@
(5.6) $\omega_{\tau_{B}\otimes\tau_{D}}^{\rm min}(F\otimes G)=(\Sigma_{[23]})^{*}(\omega_{\tau_{B}}^{\rm min}(F)\otimes\omega_{\tau_{D}}^{\rm min}(G)).$
@@
Let $A$ and $B$ be $C^{*}$-algebras, and let $\tau$ be a faithful trace on $B$. For any $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$ we have that $F$ is completely positive if and only if $\omega_{\tau}^{\rm max}(F)$ is positive.
@@
Suppose first that $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$ is completely positive. Since $\omega_{\tau}^{\rm max}(F)$ and $\mu_{\tau}^{\rm max}\circ(F\otimes\operatorname{id}_{B^{\rm op}})$ agree on the dense subspace $A\odot B^{\rm op}$, we have that $\omega_{\tau}^{\rm max}(F)$ is given by the following composition:
@@
$(\operatorname{Tr}_{n}\otimes\tau)((\operatorname{id}_{n}\otimes F)(X)Y)$ $=(\omega_{\operatorname{Tr}_{n}\otimes\tau}(\operatorname{id}_{n}\otimes F))(X\otimes Y^{\rm op})$ by (5.1)
@@
$=((\Sigma_{[23]})^{*}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(F)))(X\otimes Y^{\rm op})$ by (5.5)
@@
Note that in theorem 5.12, faithfulness of the trace is not necessary for the forward implications. We will however mostly be concerned with faithful traces in this article, so we have chosen to state the theorem in its present form.
@@
We note that if $B$ is unital and $\tau$ is a tracial state, then $\operatorname{UCP}(A,B)$ is contained in $\operatorname{TC}_{\tau}(A,B)$.
@@
Let $A$ and $B$ be $C^{*}$-algebras with faithful traces $\tau_{A}$ and $\tau_{B}$, respectively. We say that a linear map $F\colon A\to B$ is $(\tau_{A},\tau_{B})$-adjointable if there exists a linear map $F^{\sharp}\colon B\to A$ such that
(5.7) $\tau_{B}(F(a)b)=\tau_{A}(aF^{\sharp}(b))$
@@
Remark 5.17.
The requirement that $\tau_{A}$ and $\tau_{B}$ are faithful ensures that $F^{\sharp}$ is uniquely determined by $F$ through (5.7). Moreover, $F^{\sharp}$ is $(\tau_{B},\tau_{A})$-adjointable. Uniqueness of adjoints ensures that $(F^{\sharp})^{\sharp}=F$. If $C$ is a $C^{*}$-algebra with a faithful trace $\tau_{C}$ and $G$ is $(\tau_{B},\tau_{C})$-adjointable, then $G\circ F$ is $(\tau_{A},\tau_{C})$-adjointable, and uniqueness of the adjoint also implies that
(5.8) $(G\circ F)^{\sharp}=F^{\sharp}\circ G^{\sharp}.$
@@
Let $A$ and $B$ be unital $C^{*}$-algebras with faithful traces $\tau_{A}$ and $\tau_{B}$, respectively. If $F\in\operatorname{TC}_{\tau_{B}}(A,B)$ is $(\tau_{A},\tau_{B})$-adjointable, then $F^{\sharp}\in\operatorname{TC}_{\tau_{A}}(B,A)$, and we furthermore have that
(5.9) $\omega_{\tau_{A}}(F^{\sharp})=\omega_{\tau_{B}}(F)\circ\Sigma^{\rm op}.$
@@
5.2. Relation to the Choi-Jamiołkowski isomorphisms
@@
Let $n,m\in\mathbb{N}$, $A=M_{n}(\mathbb{C})$, $B=M_{m}(\mathbb{C})$, and $\tau$ be the canonical trace on $B$. Let $F\colon A\to B$ be a linear map. We recall, see for instance [Stø13, Chapter 4], that the Choi matrix $C_{F}$ of $F$ is given by
$C_{F}:=\sum_{i,j=1}^{n}e_{i,j}^{(n)}\otimes F(e_{i,j}^{(n)})\,\in\,M_{n}(\mathbb{C})\otimes M_{m}(\mathbb{C})\simeq M_{nm}(\mathbb{C}).$
The assignment $F\mapsto C_{F}$ establishes an isomorphism between $\operatorname{Hom}(A,B)$ and $A\otimes B$. This isomorphism is called the Choi-Jamiołkowski isomorphism.
@@
for all $a\in M_{n}(\mathbb{C})$, $b\in M_{m}(\mathbb{C})$, and that $\widehat{F}$ is positive if (and only if) $F$ is completely positive, in which case the positive density matrix corresponding to $\widehat{F}$ is the transpose of $C_{F}$, cf. [Stø13, Lemma 4.2.3 and Theorem 4.2.7]. Now, let $t$ denote the *-isomorphism from $B^{\rm op}=M_{m}(\mathbb{C})^{\rm op}$ onto $B=M_{m}(\mathbb{C})$ given by $t(b^{\rm op})=b^{t}$ for all $b\in B$. Then one readily verifies that
@@
Thus, if $F$ is completely positive and we identify $A\otimes B^{\rm op}$ with $A\otimes B$ via the *-isomorphism $\operatorname{id}_{A}\otimes t$, we get that the density matrix corresponding to the positive functional $\omega_{\tau}(F)$ is the positive matrix $(C_{F})^{t}$.
@@
$\widehat{F}(a\otimes e_{i,j}^{(m)})=F(a)_{i,j}\quad\text{ for every $a\in A$ and $i,j\in\{1,\ldots,m\}$},$
where $F(a)_{i,j}$ denotes the $(i,j)$-coefficient of the matrix $F(a)$. Then, as shown in [BO08, Proposition 1.5.14], the map $F$ is completely positive if and only if $\widehat{F}$ is positive, and the map $F\mapsto\widehat{F}$ gives a bijective correspondence between $\operatorname{CP}(A,M_{m}(\mathbb{C}))$ and the cone of all positive linear functionals on $A\otimes M_{m}(\mathbb{C})$. Now, it is easy to verify that $\widehat{F}(a\otimes b)={\rm Tr}_{m}(F(a)b^{t})$ for all $a\in A$ and $b\in M_{m}(\mathbb{C})$. Thus, in analogy with example 5.19, $\widehat{F}$ might be called the dual functional of $F$ on $A\otimes M_{m}(\mathbb{C})$. Moreover, setting $t(b^{\rm op}):=b^{t}$ for all $b\in M_{m}(\mathbb{C})$, it is clear that we have $\omega_{\tau}(F)=\widehat{F}\circ({\rm id}_{A}\otimes t)$.
@@
In a von Neumann algebraic setting, there exist some infinite-dimensional versions of the Choi-Jamiołkowski isomorphism in the literature. See for example [Haa21] and [HKS24] and references therein.
Remark 5.23.
When $A$ and $B$ are von Neumann algebras equipped with faithful normal states $\mu$ and $\nu$, respectively, Duvenhage and Snyman introduce in [DS18, Definition 2.10] a notion of a coupling $\omega$ of $(A,\mu)$ and $(B,\nu)$, which is a certain state on $A\odot B^{\prime}$, where $B^{\prime}$ denotes the commutant of $B$. Moreover, they associate to each linear map $E:A\to B$ a linear functional $\omega_{E}$ on $A\odot B^{\prime}$ and show in [DS18, Proposition 3.9] that $\omega_{E}$ is a coupling of $(A,\mu)$ and $(B,\nu)$ if and only if $E\in\operatorname{UCP}(A,B)$ and $\nu\circ E=\mu$. In the case where $\nu$ is tracial, it would be interesting to know if there is a connection between their result and our version of the Choi-Jamiołkowski map.
6. Embedding-induced metrics
In light of the results of section 5, we are able to define the embedding-induced metrics alluded to in (3.5).
Definition 6.1.
Let $A$ be a unital $C^{*}$-algebra and $B$ be a $C^{*}$-algebra with a faithful trace $\tau$. Let $L\colon A\otimes_{\rm max}B^{\rm op}\to[0,\infty]$ be a seminorm on $A\otimes_{\rm max}B^{\rm op}$. Define the extended metric $\Delta_{\tau,L}^{\rm max}\colon\operatorname{TC}_{\tau}(A,B)\times\operatorname{TC}_{\tau}(A,B)\to[0,\infty]$ by setting
(6.1) $\Delta_{\tau,L}^{\rm max}(F,G):=\operatorname{mk}_{L}(\omega_{\tau}^{\rm max}(F),\omega_{\tau}^{\rm max}(G)).$
In the case where $B$ is unital, $\tau$ is amenable, and $L\colon A\otimes_{\rm min}B^{\rm op}\to[0,\infty]$ is a seminorm on $A\otimes_{\rm min}B^{\rm op}$, we define $\Delta_{\tau,L}^{\rm min}\colon\operatorname{TC}_{\tau}(A,B)\times\operatorname{TC}_{\tau}(A,B)\to[0,\infty]$ by setting
(6.2) $\Delta_{\tau,L}^{\rm min}(F,G):=\operatorname{mk}_{L}(\omega_{\tau}^{\rm min}(F),\omega_{\tau}^{\rm min}(G)).$
An extended metric induced in this way will be called an embedding-induced extended metric from a seminorm.
We will determine some mild criteria under which we can guarantee that stability and chaining hold for these extended metrics. The main result for stability is theorem 6.4, while the main results for chaining are theorem 6.11, and corollary 6.13.Note that, like with the pullback-induced metrics discussed in section 4, it is generally not obvious when $\Delta_{\tau,L}^{\rm max}$ is a bounded metric on $\operatorname{TC}_{\tau}(A,B)$ for $C^{*}$-algebras $A$ and $B$. However, analogously to the situation in section 4, $\Delta_{\tau,L}^{\rm max}$ is a bounded metric when $(A\otimes_{\rm max}B^{\rm op},L)$ is a compact quantum metric space. Note that a necessary condition for this to be the case is that both $A$ and $B$ are unital. The corresponding statement holds true for $\Delta_{\tau,L}^{\rm min}$ and $(A\otimes_{\rm min}B^{\rm op},L)$.
6.1. Stability